Consider {0,2,4} as a subset of Z6. show it is a subring and does it have a unity?


1
Expert's answer
2021-09-17T03:45:26-0400

Let us consider "R=\{ 0,2,4\}" as a subset of "\Z_6". Let us show that "R" is a subring of "\Z_6." Since "0-0=0\in\Z_6,\ 2-0=2\in\Z_6,\ 4-0=4\in\Z_6,\ 0-2=4\in\Z_6,\ 2-2=0\in\Z_6,\ 4-2=2\in\Z_6,\ 0-4=2\in\Z_6,\ 2-4=4\in\Z_6,\ 4-4=0\in\Z_6,\"

we conclude that the operation of substraction is closed on "R."

Taking into account that "0\cdot 0=0\cdot 2=2\cdot 0=4\cdot 0=0\cdot 4=0\in\Z_6,\ 2\cdot 2=4\in\Z_6, \ 2\cdot 4=4\cdot 2=2\in\Z_6" and "4\cdot 4=4\in\Z_6," we conclude that the operation of multiplication is closed on "R."

Therefore, "R" is a subring of "\Z_6."


Taking into account that "4\cdot 4=4, 4\cdot 0=0\cdot 4=0,4\cdot 2=2\cdot 4=2," we conclude that "4" is unity of "R."